# {textblock*}

### v2

@@ IV Numerical results
However, Case 3 is excluded at more than the $2\sigma$ level based on the P-ACT-LB constraint on $\alpha_{s}$. This is due to the fact that, in our model, the production of PBHs with masses $\gtrsim 10^{20}\mathrm{g}$ is generally disfavored by ACT, as it is associated with a relatively large running of the spectral index (see Fig. 3). As a result, we conclude that PBHs larger than $\sim 10^{20}$ g are excluded as a dark matter candidate in our model at 2 $\sigma$ level according to the P-ACT results.
@@
The full power spectra resulting from parameter Case 2 with varying $\xi$s are shown in Fig. 4. The variation of the spectral index and its running can be observed in Fig. 4, and more intuitively in Fig. 5 and Fig. 6, where the tendencies are consistent with the analytical approximations given by Eq. (14).
@@
It is also worth mentioning that, as illustrated in Fig. 4, the enhancement of the small-scale power spectrum not only affects the large-scale spectral index $n_{s}$ and its running $\alpha_{s}$, which can be probed through direct measurements of the primordial power spectrum from future CMB surveys [95, 96], but also induces significant $\mu$-distortion [97, 98]—potentially detectable by PIXIE [29, 28] [^6] See detailed analysis in [11].. Additionally, the associated induced gravitational waves fall within the sensitivity ranges of future detectors such as LISA [31], DECIGO [100], Taiji [101]/TianQin [102] and BBO [71]. Thus, upcoming CMB and gravitational wave missions may provide critical observational tests of this scenario.
@@
Figure 5: The $n_{s}-r$ relation in the $R^{2}$ model and $\chi$-extended $R^{2}$ model shown alongside the joint constraints from BICEP/Keck observations [103] (denoted as BK18) with Planck-LB (orange), ACT-LB (blue), and P-ACT-LB (purple). The constraint contours are taken from [6]. The dark blue dashed line with hollow dots refers to the $n_{s}-r$ relation in the $R^{2}$ model, and the light cyan dashed line and markers represent the $\chi$-extended $R^{2}$ model with PBH production. For the light cyan plot, we use the model parameters from Table 1 Case 2 except for a varying $\xi$. From lower $n_{s}$ (left) to higher $n_{s}$ (right), the coupling constants used are $\xi=0.3125$ (hollow dot), 0.07250.07250.0725, 0.07150.07150.0715, 0.07050.07050.0705, 0.06950.06950.0695, 0.06890.06890.0689 (star-shaped, Case 2), 0.06850.06850.0685.
@@
Figure 6: The $n_{s}-\alpha_{s}$ relation in $\chi$-extended $R^{2}$ model shown alongside the joint constraints from Planck, P-ACT and P-ACT-LB (The constraint contours are taken from [6]). The light cyan dots refer to the $\alpha_{s}-n_{s}$ relation in $\chi$-extended $R^{2}$ model with PBH production, for which we use the model parameters from Table 1 Case 2 except for a varying $\xi$. From smaller $n_{s}$ (left) to larger $n_{s}$ (right), we take $\xi=0.3125$ (hollow dot), 0.07250.07250.0725, 0.07150.07150.0715, 0.07050.07050.0705, 0.06950.06950.0695, 0.06890.06890.0689 (star-shaped, Case 2), 0.06850.06850.0685.
V Discussion and Conclusion
@@
The recently released ACT data, combined with Planck observations, report a slightly larger scalar spectral index $n_{s}$ and a positive running $\alpha$ placing well-motivated $R^{2}$ inflation under tension at $2\sigma$ level. In this paper, we have shown that the tension can be resolved by introducing a non-minimally coupled scalar field $\chi$ to the $R^{2}$ model. By tuning the non-minimal coupling constant $\xi$, the model realizes a hybrid-like inflationary scenario, in which the non-minimally coupled scalar field $\chi$ contributes a blue-tilted component to the primordial power spectrum, yielding an increase in the $n_{s}$ and a small but positive $\alpha_{s}$, both consistent with current observational constraints.
@@
At the same time, the blue-tilted power spectrum significantly enhances the fluctuation amplitude at small scales, potentially leading to the formation of primordial black holes (PBHs) that could constitute the entire dark matter abundance. In Fig. 3, we illustrate the relationship between the spectral index $n_{s}$, its running $\alpha_{s}$, and the resulting PBH mass under the assumption that PBHs account for all dark matter.
@@
Finally, we should stress the intriguing connection between the inflation scale and the small neutrino masses $m_{\nu_{i}}$. The $\chi$ boson has coupling to the right handed neutrinos as $g_{i}\chi N_{Ri}N_{Ri}$ which generates the large Majorana masses $M_{i}$ for the right-handed neutrinos after the total inflation ends as explained in the introduction. Through the seesaw mechanism [38, 39, 40, 41], the observed neutrino masses $m_{\nu_{i}}$ is given by $m_{\nu_{i}}\simeq h_{i}^{2}\langle H\rangle^{2}/M_{i}$. Here, the $H$ is the standard-model Higgs boson, $M_{i}\simeq g_{i}\langle\chi\rangle$, and the $\langle\chi\rangle=v$ is determined by the inflation scale as shown in the text. Therefore, it is very interesting that there is a strong correlation between the inflation scale and the observed small neutrino masses.
Acknowledgements.

### v3

@@ 4 Numerical results
However, Case 3 is excluded at more than the $2\sigma$ level based on the P-ACT-LB constraint on $\alpha_{s}$. This is due to the fact that, in our model, the production of PBHs with masses $\gtrsim 10^{20}\mathrm{g}$ is generally disfavored by ACT, as it is associated with a relatively large running of the spectral index (see figure 3). As a result, we conclude that PBHs larger than $\sim 10^{20}$ g are excluded as a dark matter candidate in our model at 2 $\sigma$ level according to the P-ACT results.
@@
[…]
@@
It is also worth mentioning that, as illustrated in figure 4, the enhancement of the small-scale power spectrum not only affects the large-scale spectral index $n_{s}$ and its running $\alpha_{s}$, which can be probed through direct measurements of the primordial power spectrum from future CMB surveys [1, 6], but also induces significant $\mu$-distortion [27, 26]—potentially detectable by PIXIE [2, 74] [^5] See detailed analysis in [71]..
Additionally, the associated induced gravitational waves fall within the sensitivity ranges of future detectors such as LISA [10], DECIGO [68], Taiji [60]/TianQin [80] and BBO [51]. Thus, upcoming CMB and gravitational wave missions may provide critical observational tests of this scenario.
@@
Figure 5: The $n_{s}-r$ relation in the $R^{2}$ model and $\chi$-extended $R^{2}$ model shown alongside the joint constraints from BICEP/Keck observations [5] (denoted as BK18) with Planck-LB (orange), ACT-LB (blue), and P-ACT-LB (purple). The constraint contours are taken from [17]. The dark blue dashed line with hollow dots refers to the $n_{s}-r$ relation in the $R^{2}$ model, and the light cyan dashed line and markers represent the $\chi$-extended $R^{2}$ model with PBH production. For the light cyan plot, we use the model parameters from Table 1, Case 2 except for a varying $\xi$. From lower $n_{s}$ (left) to higher $n_{s}$ (right), the coupling constants used are $\xi=0.3125$ (hollow dot), $0.0725$, $0.0715$, $0.0705$, $0.0695$, $0.0689$ (star-shaped, Case 2), $0.0685$.
@@
Figure 6: The $n_{s}-\alpha_{s}$ relation in $\chi$-extended $R^{2}$ model shown alongside the joint constraints from Planck, P-ACT and P-ACT-LB (The constraint contours are taken from [17]). The light cyan dots refer to the $\alpha_{s}-n_{s}$ relation in $\chi$-extended $R^{2}$ model with PBH production, for which we use the model parameters from Table 1 Case 2 except for a varying $\xi$. From smaller $n_{s}$ (left) to larger $n_{s}$ (right), we take $\xi=0.3125$ (hollow dot), $0.0725$, $0.0715$, $0.0705$, $0.0695$, $0.0689$ (star-shaped, Case 2), $0.0685$.
5 Discussion and conclusion
@@
The recently released ACT data, combined with Planck observations, report a slightly larger scalar spectral index $n_{s}$ and a positive running $\alpha_{s}$ placing well-motivated $R^{2}$ inflation under tension at $2\sigma$ level. In this paper, we have shown that the tension can be resolved by introducing a non-minimally coupled scalar field $\chi$ to the $R^{2}$ model. By tuning the non-minimal coupling constant $\xi$, the model realizes a hybrid-like inflationary scenario, in which the non-minimally coupled scalar field $\chi$ contributes a blue-tilted component to the primordial power spectrum, yielding an increase in the $n_{s}$ and a small but positive $\alpha_{s}$, both consistent with current observational constraints.
@@
At the same time, the blue-tilted power spectrum significantly enhances the fluctuation amplitude at small scales, potentially leading to the formation of primordial black holes (PBHs) that could constitute the entire dark matter abundance. In figure 3, we illustrate the relationship between the spectral index $n_{s}$, its running $\alpha_{s}$, and the resulting PBH mass under the assumption that PBHs account for all dark matter.
@@
Finally, we should stress the intriguing connection between the inflation scale and the small neutrino masses $m_{\nu_{i}}$. The $\chi$ boson has coupling to the right handed neutrinos as $g_{i}\chi N_{Ri}N_{Ri}$ which generates the large Majorana masses $M_{i}$ for the right-handed neutrinos after the total inflation ends as explained in the introduction. Through the seesaw mechanism [86, 112, 113, 44], the observed neutrino masses $m_{\nu_{i}}$ are given by $m_{\nu_{i}}\simeq h_{i}^{2}\langle H\rangle^{2}/M_{i}$. Here, the $H$ is the standard-model Higgs boson, $M_{i}\simeq g_{i}\langle\chi\rangle$, and the $\langle\chi\rangle=v$ is determined by the inflation scale as shown in the text. Therefore, it is very interesting that there is a strong correlation between the inflation scale and the observed small neutrino masses.
Acknowledgments